OpenStudy (anonymous):

Find dy/dx by implicit differentiation: x^4(x+y)=y^2(3x-y)

OpenStudy (anonymous):

take the derivative of both sides with respect to x, where y is the function y(x)

OpenStudy (anonymous):

Im a little confused as to, do I chain rule it? Or product rule it?

OpenStudy (anonymous):

You can use both rules to solve it. x^5+x^4*y=3x*y^2-y^3 To take the derivative of say 3 x y^2, first use the product rule: d/dx (3x y^2)=3*y^2+3x*d/dx (y^2). Next use the chain rule: d/dx (y(x)^2) = 2y(x)*y'(x) Putting it all together: d/dx (3x y^2)=3*y^2+3x*2y*y'.